# Multi-Modal Multi-Task (M3T) Federated Foundation Models for Embodied AI: Potentials and Challenges for Edge Integration

### v1

@@ VI Conclusion
In this paper, we envisioned that by integrating the generalization power of M3T FMs with the decentralized learning capabilities of FL, FFMs can offer a unified learning approach in the embodied AI ecosystems. Through the EMBODY framework, we articulated the core dimensions that should be addressed for FFMs to succeed in real-world embodied AI deployments. For each dimension, we identified open challenges and envisioned actionable research directions that span model architecture, training dynamics, and system configuration. We also discussed evaluation protocols and trade-off analysis for FFM deployment over embodied AI.

### v2

@@ VII Conclusion
In this paper, we envisioned that by integrating the generalization power of M3T-FMs with the decentralized learning capabilities of FL, M3T-FFMs can offer a unified learning approach in the embodied AI ecosystems. Through the EMBODY framework, we articulated the core dimensions that should be addressed for M3T-FFMs to succeed in real-world embodied AI deployments. For each dimension, we identified open challenges and envisioned actionable research directions that span model architecture, training dynamics, and system configuration. We also discussed evaluation protocols and trade-off analysis for M3T-FFM deployment over embodied AI. We highlight that EMBODY provides a structured foundation that can serve as a benchmark framework for the community. By aligning future work with the EMBODY dimensions and their associated evaluation metrics, researchers can systematically compare methods, quantify trade-offs, and ensure that research progress is both measurable and reproducible.